OpenStudy (anonymous):

find the exact value of 2 cos^2 (pi/8) -1

OpenStudy (anonymous):

Trig identities can get this simpler. Did my last response show up? I've got bad connection at the moment.

OpenStudy (anonymous):

didn't show up . im still pretty confused

OpenStudy (anonymous):

\[\cos ^{2} (x)+\sin^{2}(x)=1\] Double both sides: \[2\cos ^{2} (x)+2\sin^{2}(x)=2\] Subtract 1 from each side: \[2\cos ^{2} (x)+2\sin^{2}(x)-1=1\] Now do you see your equation?